I have a problem remembering how to perform the following equation. Please advise.
(-1/2)^-3

Answer by nyc_function(2733)

What equation?
You cannot have an equation without an equal sign.
(-1/2)^-3
Use the rule:
(a)^(negative exponent) = 1/(a)^(positive exponent), where "a" is any integer.
(-1/2)^-3 becomes 1/(-1/2)^3 = 1/(-1/8) = 8/-1 = -8